# captionUnknown document class (or package)

### v1

@@ 8 Conclusion
This survey paper provides a thorough exploration of the integration of LLMs with 3D data. Systematically reviewing methodologies, applications, and emergent abilities of LLMs in processing, understanding, and generating 3D data, the survey underlines the transformative potential of LLMs across a spectrum of 3D tasks. From enhancing spatial comprehension and interaction within 3D environments to driving forward the capabilities of embodied AI systems, LLMs emerge as pivotal in advancing the field.
Key findings include the identification of LLMs’ unique advantages such as zero-shot learning, advanced reasoning, and extensive world knowledge, which are instrumental in bridging the gap between textual information and spatial interpretation. The paper showcases the wide array of tasks where LLMs’ integration with 3D data has been successfully demonstrated. The exploration of other 3D vision-language methods alongside LLMs reveals a rich landscape of research aiming to deepen our understanding of the 3D world.
Furthermore, the survey highlights significant challenges such as data representation, model scalability, and computational efficiency, suggesting that overcoming these hurdles is crucial for the full realization of LLMs’ potential in 3D applications. In conclusion, this survey not only offers a comprehensive overview of the current state of 3D tasks using LLMs but also sets the stage for future research directions. It calls for a collaborative effort to explore and expand the capabilities of LLMs in understanding and interacting with the complex 3D world, paving the way for further advancements in the area of spatial intelligence.

### v2

@@ 8 Conclusion
This survey paper provides a thorough exploration of the integration of LLMs with 3D data. Systematically reviewing methodologies, applications, and emergent abilities of LLMs in processing, understanding, and generating 3D data, the survey underlines the transformative potential of LLMs across a spectrum of 3D tasks. Key findings include the identification of LLMs’ unique advantages such as zero-shot learning, advanced reasoning, and extensive world knowledge, which are instrumental in bridging the gap between textual information and spatial interpretation. The paper showcases the wide array of tasks where LLMs’ integration with 3D data has been successfully demonstrated. The exploration of other 3D vision-language methods alongside LLMs reveals a rich landscape of research aiming to deepen our understanding of the 3D world.
Furthermore, the survey highlights significant challenges such as data representation, model scalability, acquisition of spatial reasoning, data bottleneck and evaluation metrics, suggesting that overcoming these hurdles is crucial for the full realization of LLMs’ potential in 3D applications. In conclusion, this survey not only offers a comprehensive overview of the current state of 3D tasks using LLMs but also sets the stage for future research directions. It calls for a collaborative effort to explore and expand the capabilities of LLMs in understanding and interacting with the complex 3D world, paving the way for further advancements in the area of spatial intelligence.